OpenStudy (anonymous):

how would you solve (5xy^3=5)^2=x^2

OpenStudy (anonymous):

That is typed exactly as it is shown in your book?

OpenStudy (anonymous):

oh sorry, it's (5xy^3+5)^2=x^2

OpenStudy (anonymous):

What do you have to solve for? There is only 1 equation and two unknowns so you won't be able to get a numerical answer

OpenStudy (anonymous):

\[(5xy^3+5)^2=x^2\]or\[(5xy^{3+5})^2=x^2\]first or second one?

OpenStudy (anonymous):

you have to use implicit differentiation to find the "y" prime in terms of x and y; the first one ( :

OpenStudy (anonymous):

The first thing I would do is get rid of the power of two by taking the square root of both sides. Makes it easier

OpenStudy (anonymous):

\[5xy^3+5=x\]

OpenStudy (anonymous):

\[{d \over dx}[5xy^3+5=x]=5x2y^2y'+5y^3 + 0 = 1\]You have to do the product rule for the 5xy^3 term. Everytime you take the derivative of y you have to out a y' with it. The last steo is to just solve for y' algebraically

OpenStudy (anonymous):

Did that make sense?

OpenStudy (anonymous):

yes, that made sense. thank you so much . you saved

OpenStudy (anonymous):

Don't forget to simplify, the y;s should be able to reduce

OpenStudy (anonymous):

yes, yes i got it (: